OpenStudy (anonymous):

How many zeros will be in the standard form of the exponential expression ten to the ninety-ninth power? ten zeros ninety-eight zeros ninety-nine zeros one hundred zeros

OpenStudy (anonymous):

10.0x10^99=10+99 zeros 100 zeros

zepdrix (zepdrix):

\[\Large 10^1\quad=\quad 10\qquad\qquad\text{(one zero)}\]\[\Large 10^2\quad=\quad 100\qquad\;\quad \text{(two zeros)}\]\[\Large ...\]\[\Large 10^{99}\quad=\quad ... \qquad\;\;\quad \text{how many zeros?}\]